# Lipid membrane domains control actin network viscoelasticity

### v1

@@ Abstract
preprint: APS/123-QED
@@
Lipid bilayer membranes are 2D interfacial materials that elastically resist bending and stretching, but exhibit viscous in-plane behavior Israelachvili (2011). In the mammalian cell, a network of semiflexible actin filaments underlies and applies forces to the plasma membrane, allowing the cell mechanically interact with its environment Alberts et al. (2017); Phillips et al. (2009). In addition to generating forces orthogonal to the membrane, the cytoskeleton also plays a role in organizing the spatial distribution and movement of protein and lipid clusters on the cell surface Kusumi and Sako (1996, 1996); Fujiwara et al. (2016); Viola and Gupta (2007).
@@
Reconstituted lipid membranes are often used to study the interactions between heterogeneous membranes and the viscoelastic actin cytoskeleton Arnold et al. (2023); Arnold and Takatori (2023); Gubbala et al. (2024); Honigmann et al. (2014); Vogel et al. (2017). Multicomponent lipid membranes can undergo 2D liquid-liquid phase separation at room temperature Veatch and Keller (2003), and lipid compositions can be tuned to make actin selectively adsorb onto a single phase of the membrane Arnold et al. (2023); Arnold and Takatori (2023); Honigmann et al. (2014); Vogel et al. (2017); Schroer et al. (2020). A dense 2D concentration of filamentous actin can form viscoelastic networks on the membrane that alter the lipid phase behavior Arnold and Takatori (2023), much like 3D elastic polymer gels, which alter the growth of embedded liquid inclusions Style et al. (2018); Rosowski et al. (2020a, b); Fernández-Rico et al. (2022, 2023); Meng et al. (2024); Vidal-Henriquez and Zwicker (2021); Wei et al. (2020). However, unlike prior 3D solid/liquid composite materials whose steady droplets are much larger than the characteristic mesh size of the polymer network Style et al. (2018); Rosowski et al. (2020a, b); Fernández-Rico et al. (2022, 2023); Meng et al. (2024); Vidal-Henriquez and Zwicker (2021); Wei et al. (2020), the mesh size of the actin networks are typically 1-10 µm, comparable to the lipid domain size on reconstituted membranes. Thus, unlike prior studies on 3D solid/liquid composites, the 2D lipid domains do not experience continuum-scale interactions with the actin network, and the domains instead adopt highly irregular shapes as they conform to the network topology Arnold et al. (2023); Arnold and Takatori (2023); Gubbala et al. (2024); Honigmann et al. (2014); Vogel et al. (2017). Moreover, due to the two-dimensionality of the system, even thin ($\approx$7 nm) actin filaments form impenetrable barriers that confine the domains without access to the third dimension Arnold and Takatori (2023).
@@
In this article, we present a composite 2D material in which liquid droplets control the viscoelasticity of an adsorbed actin network. Elastic actin bundles are adsorbed to the liquid-disordered (Ld) continuous phase of a two-phase membrane without wetting liquid-ordered (Lo) dispersed domains (Fig. 1A). We recently showed that that active myosin-driven flows accelerate domain coarsening by 2$\times$ compared to thermal coarsening in such materials Arnold et al. (2023). Here, we consider the two-way coupling between actin viscoelasticity and the Lo domain morphology and structure in the absence of myosin-driven activity. We find that the rigid actin bundles constrain the domains to adopt angular morphologies, deforming the Lo/Ld interface against line tension. The domains, in turn, accelerate the viscoelastic relaxation of the actin network, decreasing the relaxation time $\tau$ by $\approx$2.5$\times$. Finally, we rapidly change the chemical composition of the membrane, finding that the network stiffens as the domains grow, which provides a tool for controlling the mechanical properties of 2D composite interfaces.
@@
Results
Membrane composite material design
@@ Results
We achieve a planar, phase-separated lipid bilayer by rupturing a giant unilamellar vesicle (GUV) containing 46% 1,2-dioleoyl-sn-glycero-3-phosphocholine (DOPC), 35% dipalmitoylphosphatidylcholine (DPPC), 15% cholesterol, 4% 1,2-dioleoyl-3-trimethylammonium propane (DOTAP), and trace 1,2-distearoyl-sn-glycero-3-phosphoethanolamine-N-[poly(ethylene glycol)2000-N’-carboxyfluorescein] (DSPE-PEG2k-FITC, Lo-partitioning fluorescent dye) onto a polymer-cushioned cover slip (see Methods), causing the membrane to spread on the surface (supplemental Fig. S1). We adsorb filamentous actin (F-actin) to the membrane at 37°C via electrostatic interactions between negatively-charged actin and positively-charged DOTAP Schroer et al. (2020); Heath et al. (2013). Upon cooling to room temperature, the membrane phase-separates, leaving Lo domains dispersed in a continuous Ld phase Veatch and Keller (2003). Once the Lo domains nucleate amidst the actin network, they completely de-wet from actin, excluding it into the Ld phase (Fig. 1A-B, supplemental Fig. S1).
@@
The fluorescence micrographs in Fig. 1C suggest that the thin $\sim\mathcal{O}$(nm) actin filaments align and cluster to form thick $\sim\mathcal{O}$(µm) bundles, as has previously been observed in electrostatically-adsorbed actin on lipid membranes Heath et al. (2013). We observe that when small amounts of actin are adsorbed to the membrane surface, the thin bundles assemble into a disordered network that avoids and deforms around the domains (Fig. 1C, left panel). However, if more actin is allowed to adsorb, the bundles thicken and ultimately span nearly the entire Ld phase (Fig. 1C, center and right panels). These thicker bundles interact with the Lo/Ld interface, straightening the domain edges next to and along the bundles, while sharply deforming the domain edges near the intersections of actin bundles (Fig. 1B-C).
Actin elasticity heterogeneously deforms domains
@@
We hypothesize that the mechanism by which actin deforms the Lo/Ld domain interface relies on actin’s bending rigidity and its tendency to nematically align on the lipid membrane. Heath et al. showed that actin filaments adsorbed electrostatically to Ld-phase supported lipid bilayers will align into a liquid crystalline structure with quasi-long ranged nematic order and periodic filament spacing Heath et al. (2013). We polymerize globular actin (G-actin) into F-actin ex situ, and then adsorb F-actin to a heated, single-phase membrane. Thus, filaments bind in all orientations, and then rotate to form parallel bundles on the fluid surface, subject to limitations due to entanglements and crossovers between filaments.
@@
Previously, micron-sized liquid inclusions were grown in 3D composite materials with nanometer-scale spacing between elastic polymers Style et al. (2015, 2018); Rosowski et al. (2020a, b); Fernández-Rico et al. (2023). Thus, the droplets experienced effectively an isotropic elastic stress exerted by a continuum polymer gel. In our 2D material, the actin bundles and Lo domains are both $\sim\mathcal{O}$(µm) in size, causing the domains to deform anisotropically in response to actin elasticity. To characterize the anisotropic elastic stresses on domains, we fit parametric curves to the domain boundaries, find the unit normal $\mathbf{n}$, and calculate the local curvature $\nabla\cdot\mathbf{n}$ along the boundary (Fig. 2D). At low actin densities, actin exerts little influence on domain shape, resulting in relatively uniform curvature (Fig. 2D, panels 1 and 2). Meanwhile, dense actin bundles elicit large heterogeneity in curvature within each domain boundary, with high curvature concentrating in only a few pointed regions (Fig. 2D, panels 3 and 4).
@@
The surface pressures in the Ld ($P_{\mathrm{Ld}}$) and Lo ($P_{\mathrm{Lo}}$) phases have units of force/length, and are balanced by line tension $\lambda$ and a compressive elastic stress due to actin $\sigma_{\mathrm{el}}$. While equations of this form are often used in 3D composites Meng et al. (2024); Wei et al. (2020); Vidal-Henriquez and Zwicker (2021), our Lo domains do not contact actin directly, but rather experience elastic stress transmitted through the Ld phase. Thus, we combine $\sigma_{\mathrm{el}}$ and $P_{\mathrm{out}}$ into an effective outside pressure, $P_{\mathrm{Ld}}^{\mathrm{eff}}=\sigma_{\mathrm{el}}+P_{\mathrm{out}}$, which simplifies Eq. 2: $P_{\mathrm{Lo}}-P_{\mathrm{Ld}}^{\mathrm{eff}}=\lambda\nabla\cdot\mathbf{n}$.
In a quasi-static membrane, lipid flows within the purely viscous Lo phase are assumed to be minimal, suggesting that $P_{\mathrm{Lo}}$ is spatially invariant within each Lo domain. Thus, the spatial variations in $\nabla\cdot\mathbf{n}$ seen in Fig. 2D are balanced by variations in $P_{\mathrm{Ld}}^{\mathrm{eff}}$, which arise due to the actin elasticity compressing the domain anisotropically. Previous work has focused only on heterogeneities in matrix stiffness between 3D droplets, with each droplet still experiencing isotropic compression Meng et al. (2024). We hypothesize that the two-dimensionality of our material, which encourages nematic actin bundling and thus maintains parity between the network and droplet length scales, is responsible for this unusual anisotropic stress concentration. Thus, the actin elasticity dictates the lipid domain architecture, and actin density provides a mechanism of controlling droplet morphology.
@@
Coarsening lipid domains accelerate the relaxation of viscoelastic actin networks
Actin and lipid domains are two-way coupled: actin controls the domain morphology while the domains themselves control actin viscoelasticity. While the actin network structure remains quasi-static over short time scales of seconds to minutes, the structure slowly relaxes over tens of minutes to hours. In Fig. 3A, we consider the actin network structure on both a single-phase Ld membrane (top) and a two-phase Lo/Ld membrane (bottom). We use differential dynamic microscopy (DDM) to calculate the intermediate scattering function $F(k,t)$ of actin for wave vector magnitude $k=|\mathbf{k}|$ over time $t$ (see Methods) Cerbino and Trappe (2008); Giavazzi et al. (2009). We define the mean wave vector $\langle k\rangle$ of actin in each sample using the static structure factor $S(k)$

### v2

@@ Abstract
UCSB] Department of Chemical Engineering, University of California, Santa Barbara, Santa Barbara, CA, 93106 UCSB] Department of Chemical Engineering, University of California, Santa Barbara, Santa Barbara, CA, 93106
{tocentry}
1 Introduction
@@
Lipid bilayer membranes are 2D interfacial materials that elastically resist bending and stretching, but exhibit viscous in-plane behavior 1. In the mammalian cell, a network of semiflexible actin filaments underlies and applies forces to the plasma membrane, allowing the cell to mechanically interact with its environment 2, 3. In addition to generating forces orthogonal to the membrane, the cytoskeleton also plays a role in organizing the spatial distribution and movement of protein and lipid clusters on the cell surface 4, 5, 6.
@@
Reconstituted lipid membranes are often used to study the interactions between heterogeneous membranes and the viscoelastic actin cytoskeleton 7, 8, 9, 10, 11. Multicomponent lipid membranes can undergo 2D liquid-liquid phase separation at room temperature 12, and lipid compositions can be tuned to make actin selectively adsorb onto a single phase of the membrane 7, 8, 10, 11, 13. A dense 2D concentration of filamentous actin can form viscoelastic networks on the membrane that alter the lipid phase behavior 8, much like 3D elastic polymer gels, which alter the growth of embedded liquid inclusions 14, 15, 16, 17, 18, 19, 20, 21, 22. However, unlike most prior 3D solid/liquid composite materials whose steady droplets are much larger than the characteristic mesh size of the polymer network 14, 15, 16, 17, 18, 19, 20, 21, the mesh size of the actin networks are typically 1-10 µm, comparable to the lipid domain size on reconstituted membranes. Thus, unlike prior studies on 3D solid/liquid composites, the 2D lipid domains do not experience continuum-scale interactions with the actin network, and the domains instead adopt highly irregular shapes as they conform to the network topology 7, 8, 9, 10, 11, 23. Moreover, due to the two-dimensionality of the system, even thin ($\approx$7 nm) actin filaments form impenetrable barriers that confine the domains without access to the third dimension 8.
@@
In this article, we present a composite 2D material in which liquid droplets control the viscoelasticity of an adsorbed actin network. Elastic actin bundles are adsorbed to the liquid-disordered (Ld) continuous phase of a two-phase membrane without wetting liquid-ordered (Lo) dispersed domains (Fig. 1A). We recently showed that that active myosin-driven flows accelerate domain coarsening by 2$\times$ compared to thermal coarsening in such materials 7. Here, we consider the two-way coupling between actin viscoelasticity and the Lo domain morphology and structure in the absence of myosin-driven activity. We find that the rigid actin bundles constrain the domains to adopt angular morphologies, deforming the Lo/Ld interface against line tension. The domains, in turn, accelerate the viscoelastic relaxation of the actin network, decreasing the relaxation time $\tau$ by $\approx$2.5$\times$. Finally, we rapidly change the chemical composition of the membrane, finding that the network stiffens as the domains grow, which provides a tool for controlling the mechanical properties of 2D composite interfaces.
@@
2 Experimental Methods
2.1 Buffers
[…]
2.2 Actin and heavy meromyosin preparation
Rabbit skeletal muscle actin is purified from muscle acetone powder (Pel-Freez, catalog no: 41995-2, Lot 16743) using standard methods 24, 25. No rabbits or other animals are directly involved in this study. Actin is stored as depolymerized globular actin (G-actin) at -80°C in G-buffer with 6% sucrose until use.
Heavy meromyosin (HMM) is purified from rabbit skeletal muscle (Pel-Freez Biologicals, Rogers, Arkansas) using standard methods 26. HMM is frozen in 10 mM potassium phosphate buffer (pH 7.0), 100 mM potassium chloride, 0.3 mM EGTA, 1 mM DTT, and 6% sucrose at -80°C until use.
2.3 Giant unilamellar vesicle (GUV) preparation
Giant unilamellar vesicles (GUVs) are prepared using the established electroformation method 27. Briefly, lipids are mixed with the following composition: 45.7% 1,2-dioleoyl-sn-glycero-3-phosphocholine (DOPC, Avanti catalog no: 850375P), 35% 1,2-dipalmitoyl-sn-glycero-3-phosphocholine (DPPC, Avanti catalog no: 850355C), 15% cholesterol (TCI Chemical, catalog no: C3624), 4% 1,2-dioleoyl-3-trimethylammonium-propane (DOTAP, Avanti catalog no: 890890P), and 0.3% 1,2-distearoyl-sn-glycero-3-phosphoethanolamine-N-[poly(ethylene glycol)2000-N’-carboxyfluorescein] (DSPE-PEG2k-FITC, Avanti catalog no: 810120C). This mixture of lipids and cholesterol phase-separates into liquid-ordered (Lo) domains dispersed within a liquid-disordered (Ld) continuous phase. Lipids are spread on an indium tin oxide (ITO)-coated microscope slide (Diamond Coatings, 8-12 Ohm slide) and dried under vacuum for 30 minutes.
2.4 Surface preparation
This mixture of HMM and PLL-g-PEG on the coverslip surface helps to cushion the planar lipid bilayer and create space between the membrane and glass. Lipid membrane domains experience frictional interactions with nanoscale roughness when directly supported on uncushioned glass coverslips. Thus, without the cushion of bulky HMM and PEG, lipid domains would experience kinetic arrest and fail to grow beyond the diffraction limit 28.
2.5 Adsorbing planar lipid bilayer to treated coverslip surface
2.6 Assembling actomyosin cortex on a lipid bilayer
Planar lipid bilayers adsorbed to cover slips are incubated in 1 µM F-actin for 10 minutes at 37 °C. F-actin adsorbs uniformly to the bilayer via electrostatic attraction to DOTAP 13, 29. The membrane is then cooled to room temperature, causing it to phase-separate into liquid-ordered and liquid-disordered phases. The actin then sequesters into the Ld phase. Unbound actin is washed from the cover slip with A-buffer. A complete phase-separated bilayer with actin adsorbed to the Ld phase is shown with uncropped edges in supporting Fig. S1.
2.7 Inserting cholesterol into phase-separated membrane sample
[…]
2.8 Preparing and tracking tracer actin
Every few minutes, we capture a short series of images (100 total images, 0.2 s between images) of the tracer actin filaments We track the movement of the tracer filaments using the Trackmate plugin in ImageJ 30, and calculate the variance in tracking positions over the 20-second time lapse.
2.9 Microscope for all imaging experiments
All imaging is carried out on an inverted Nikon Ti2-Eclipse microscope (Nikon Instruments) using an oil-immersion objective (Apo 100x, NA 1.45, oil). Lumencor SpectraX Multi-Line LED Light Source is used for excitation (Lumencor, Inc). Fluorescent light is spectrally filtered with emission filters (432/36, 515/30, 595/31, and 680/42; Semrock, IDEX Health and Science) and imaged on a Photometrics Prime 95 CMOS Camera (Teledyne Photometrics). Microscope images are collected using MicroManager 1.4 software 31.
2.10 Image analysis
2.10.1 Actin density
2.10.2 Domain perimeter
Domains are binarized using adaptive thresholding. In many cases, binary thresholding merges tightly-spaced domains, which will greatly increase the excess perimeter if not corrected. To separate merged domains, we take each binary domain of more than ten pixels in size and check whether it can be easily separated into multiple distinct objects. First, the binarization threshold is increased until only 30% of the original domain pixels remain. The threshold is then iteratively lowered, 2% at a time, and the number of distinct objects greater than 15 pixels in size in the resulting binary image counted. If, after any threshold decrease, the number of binary objects in the image decreases, then the layer of pixels connecting the objects is deleted. This process continues until the original threshold is recovered, effectively severing a single row of pixels between the domains. The perimeter and area of each domain is then measured using the MATLAB Image Processing Toolbox. Finally, we account for any remaining merged domains by removing any significant outliers in $L_{\mathrm{ex}}$, as determined by a Grubbs test.
2.10.3 Domain curvature
2.10.4 Relaxation time
The relaxation time of the actin network is found using principles of differential dynamic microscopy (DDM), as described previously 32, 33. The image structure function $D(\mathbf{k},t)$ is calculated by taking the fast Fourier transform of differences in image intensities $I$, separated by lag time $t$:
$D(k,t)=FFT\left[|I(\mathbf{x},t)-I(\mathbf{x},0)|^{2}\right].$ (1)
[…]
$F(k,t)=1-\frac{D(k,t)-B(k)}{A(k)}$ (2)
where $A(k)$ and $B(k)$ are time-independent constants. We use the known asymptotic behavior $F(k,t\rightarrow\infty)=0$ and $F(k,t=0)=1$ to find $A(k)$ and $B(k)$, using the relations $A(k)+B(k)=D(k,t\rightarrow\infty)$ and $B(k)=D(k,t=0)$. We find averaging the first and last few data points that qualitatively display asymptotic behavior to be the most effective way to find $F(k,t)$ in our system. While $F(k,t)$ is exponential in time in most DDM experiments32, our relaxation data is markedly non-exponential, thus making a two-parameter nonlinear fit for $A(k)$ and $B(k)$ inappropriate. To avoid assuming a functional form for our relaxation data, we also define the relaxation time $\tau$ as the half life of the curve, such that $F(k,\tau)=0.5$.
2.10.5 Domain and actin size
We find the size of lipid domains and actin using the mean wave vector of the static structure factor $S(k)$. To reduce artifacts due to photo bleaching, we first apply histogram-matched bleach correction to the time-lapse in ImageJ 35. We take the 2D fast Fourier transform (FFT) of each image and take the absolute value of the Fourier coefficients to be $S(n)=|FFT(n)|$ where $n=0,1,2,...$ is the Fourier mode. We radially bin $S(n)$ to the nearest integer mode. For example, the first diagonal point $S(n=\sqrt{2})$ would be rounded down to $n=1$ and averaged with all other coefficients $S(n=1)$. Fourier modes are converted to wave vectors using the transformation $k=2\pi n/L$ where $L$ is the image width. Only square images are analyzed, so that $L_{x}=L_{y}$, and thus $k_{x}=k_{y}$ for $n_{x}=n_{y}$.
3 Results and Discussion
3.1 Membrane composite material design
@@
We achieve a planar, phase-separated lipid bilayer by rupturing a giant unilamellar vesicle (GUV) containing 46% 1,2-dioleoyl-sn-glycero-3-phosphocholine (DOPC), 35% dipalmitoylphosphatidylcholine (DPPC), 15% cholesterol, 4% 1,2-dioleoyl-3-trimethylammonium propane (DOTAP), and trace 1,2-distearoyl-sn-glycero-3-phosphoethanolamine-N-[poly(ethylene glycol)2000-N’-carboxyfluorescein] (DSPE-PEG2k-FITC, Lo-partitioning fluorescent dye) onto a polymer-cushioned cover slip (see Experimental Methods), causing the membrane to spread on the surface (supporting Fig. S1). We adsorb filamentous actin (F-actin) to the membrane at 37°C via electrostatic interactions between negatively-charged actin and positively-charged DOTAP 13, 29. Upon cooling to room temperature, the membrane phase-separates, leaving Lo domains dispersed in a continuous Ld phase 12. Once the Lo domains nucleate amidst the actin network, they completely de-wet from actin, excluding it into the Ld phase (Fig. 1A-B, supporting Fig. S1).
We previously observed this near-total partitioning of actin into the Ld phase in similar experiments 7, 8, and expect it to be caused in part by the imbalance of charge between the two phases, as DOTAP strongly partitions into the Ld phase 10. However, simulations have shown that when actin adsorbs electrostatically to a lipid bilayer, the bilayer deforms out-of-plane to wet the actin 13. Thus, we expect that the stiffer Lo phase also resists the necessary bending to wet actin, ensuring near-total partitioning into the Ld phase.
@@
The fluorescence micrographs in Fig. 1C suggest that the thin $\sim\mathcal{O}$(nm) actin filaments align and cluster to form thick $\sim\mathcal{O}$(µm) bundles, as has previously been observed in electrostatically-adsorbed actin on lipid membranes 29. We observe that when small amounts of actin are adsorbed to the membrane surface, the thin bundles assemble into a disordered network that avoids and deforms around the domains (Fig. 1C, left panel). However, if more actin is allowed to adsorb, the bundles thicken and ultimately span nearly the entire Ld phase (Fig. 1C, center and right panels). These thicker bundles interact with the Lo/Ld interface, straightening the domain edges next to and along the bundles, while sharply deforming the domain edges near the intersections of actin bundles (Fig. 1B-C).
3.2 Actin elasticity heterogeneously deforms domains
@@
We hypothesize that the mechanism by which actin deforms the Lo/Ld domain interface relies on actin’s bending rigidity and its tendency to nematically align on the lipid membrane. Heath et al. showed that actin filaments adsorbed electrostatically to Ld-phase supported lipid bilayers will align into a liquid crystalline structure with quasi-long ranged nematic order and periodic filament spacing 29. We polymerize globular actin (G-actin) into F-actin ex situ, and then adsorb F-actin to a heated, single-phase membrane. Thus, filaments bind in all orientations, and then rotate to form parallel bundles on the fluid surface, subject to limitations due to entanglements and crossovers between filaments.
@@
Previously, micron-sized liquid inclusions were grown in 3D composite materials with nanometer-scale spacing between elastic polymers 38, 14, 15, 16, 18. Thus, the droplets experienced effectively an isotropic elastic stress exerted by a continuum polymer gel. In our 2D material, the actin bundles and Lo domains are both $\sim\mathcal{O}$(µm) in size, causing the domains to deform anisotropically in response to actin elasticity. To characterize the anisotropic elastic stresses on domains, we fit parametric curves to the domain boundaries, find the unit normal $\mathbf{n}$, and calculate the local curvature $\nabla\cdot\mathbf{n}$ along the boundary (Fig. 2D). At low actin densities, actin exerts little influence on domain shape, resulting in relatively uniform curvature (Fig. 2D, panels 1 and 2). Meanwhile, dense actin bundles elicit large heterogeneity in curvature within each domain boundary, with high curvature concentrating in only a few pointed regions (Fig. 2D, panels 3 and 4).
@@
The surface pressures in the Ld ($P_{\mathrm{Ld}}$) and Lo ($P_{\mathrm{Lo}}$) phases have units of force/length, and are balanced by line tension $\lambda$ and a compressive elastic stress due to actin $\sigma_{\mathrm{el}}$. While equations of this form are often used in 3D composites 19, 21, 20, our Lo domains do not contact actin directly, but rather experience elastic stress transmitted through the Ld phase. Thus, we combine $\sigma_{\mathrm{el}}$ and $P_{\mathrm{out}}$ into an effective outside pressure, $P_{\mathrm{Ld}}^{\mathrm{eff}}=\sigma_{\mathrm{el}}+P_{\mathrm{out}}$, which simplifies Eq. 4: $P_{\mathrm{Lo}}-P_{\mathrm{Ld}}^{\mathrm{eff}}=\lambda\nabla\cdot\mathbf{n}$.
In a quasi-static membrane, lipid flows within the purely viscous Lo phase are assumed to be minimal, suggesting that $P_{\mathrm{Lo}}$ is spatially invariant within each Lo domain. Thus, the spatial variations in $\nabla\cdot\mathbf{n}$ seen in Fig. 2D are balanced by variations in $P_{\mathrm{Ld}}^{\mathrm{eff}}$, which arise due to the actin elasticity compressing the domain anisotropically. Previous work has focused only on heterogeneities in matrix stiffness between 3D droplets, with each droplet still experiencing isotropic compression 19. We hypothesize that the two-dimensionality of our material, which encourages nematic actin bundling and thus maintains parity between the network and droplet length scales, is responsible for this unusual anisotropic stress concentration. Thus, the actin elasticity dictates the lipid domain architecture, and actin density provides a mechanism of controlling droplet morphology.
@@
3.3 Coarsening lipid domains accelerate the relaxation of viscoelastic actin networks
Actin and lipid domains are two-way coupled: actin controls the domain morphology while the domains themselves control actin viscoelasticity. While the actin network structure remains quasi-static over short time scales of seconds to minutes, the structure slowly relaxes over tens of minutes to hours. In Fig. 3A, we consider the actin network structure on both a single-phase Ld membrane (top) and a two-phase Lo/Ld membrane (bottom). Due to the high variability in actin adsorption to the membrane between trials, we present a single representative pair of membranes with similar actin densities. We use differential dynamic microscopy (DDM) to calculate the intermediate scattering function $F(k,t)$ of actin for wave vector magnitude $k=|\mathbf{k}|$ over time $t$ (see Methods) 32, 33. We define the mean wave vector $\langle k\rangle$ of actin in each sample using the static structure factor $S(k)$